A racing car moves on a circular track of radius b. The car starts from rest and its speed increases at a constant rate "\alpha". Find the angle between its velocity and acceleration vectors at time t.


1
Expert's answer
2021-10-05T15:37:29-0400
"|a_{\tau}|=\alpha, v=\alpha t"

"|a_r|=\dfrac{v^2}{b}=\dfrac{\alpha^2t^2}{b}"

"|a|=\sqrt{(\alpha)^2+(\dfrac{\alpha^2t^2}{b})^2}"

"\cos\theta=\dfrac{|a_{\tau}|}{|a|}=\dfrac{\alpha}{\sqrt{(\alpha)^2+(\dfrac{\alpha^2t^2}{b})^2}}"

"=\dfrac{b}{\sqrt{b^2+\alpha^2t^4}}"

"\theta=\cos^{-1}(\dfrac{b}{\sqrt{b^2+\alpha^2t^4}})"